# A Topic-wise Exploration of the Telegram Group-verse

### v1

@@ Appendix A Ethics
In the first place, while crawling TGStat, we respect the page visit rate imposed by the website itself. We keep our data extraction rate (few pages per minute) at a level to minimise TGStat’s load. For the same reason, we repeat the crawling only once a week.
@@
Monitoring Telegram groups can also be delicate. First of all, we checked the privacy policy of the Telegram platform. It does not forbid crawling. We also asked explicitly Telegram’s support at support@telegram.org and security@telegram.org to declare our intention, asking them to share with us any restrictions or limitations. We received no answer.
@@
To respect privacy restrictions imposed by users and administrators, we restrict our analysis to public groups where it is common knowledge and well-accepted that anybody with a link to access the chat can participate as a member — if not as an active user. We respect both the groups where the administrator sets the auto-delete functionality and those where admins refuse — or ignore — our join request. We do not monitor nor store any data about these groups.
Considering users participating in groups, we store only their TelegramID (which are randomly generated), and do not store any Personally Identifiable Information such as usernames (which are in any case freely chosen by the users and can change at any time) or profile pictures. Usernames can appear in the message content when a user is mentioned by someone else. We did not make any effort to re-identify any users.
@@
Finally, we are aware that Telegram is sometimes used to share copyright-protected material and illicit content — more so, our results suggest that this behaviour is frequent. To prevent being implicated in such activity, we avoid downloading and storing any actual media such as pictures and videos, only collecting metadata (e.g., video duration and size).
Appendix B Appendix
In the appendix, we include some secondary results that can help the reader have a more complete understanding of the context of our analysis.
In Figure 7, we show the breakdown of the groups we found on TGStat, and those we actually monitor for the paper.
In Table 2, we provide a qualitative explanation of the topics under observation.
In Table 3, we present per-topic detailed quantities.
Figure 7. Per-topic groups of TGStat and results of our crawling.
@@ Appendix B Appendix
Table 3. Per-topic detailed metrics.
Topic English groups Russian groups Bot msgs (%) Avg msg length (ch.) Poll (%) Voice (%) Image (%) Emoji (%) Video (%) Sticker (%) GIF (%) Link (%) Avg video size (MB) Avg video duration (min)
Education 46 13 8.9 43.3 0.01 0.02 5.13 12.43 0.04 0.01 $<$0.01 0.06 49.9 4.3
Bookmaking 24 49 6.4 68.3 $<$0.01 0.03 8.67 19.81 0.40 0.72 0.38 0.04 14.1 0.8
Crypto 45 15 8.3 44.4 $<$0.01 $<$0.01 3.89 12.54 0.04 0.81 0.39 0.14 12.5 0.9
Technologies 23 30 9.9 58.5 $<$0.01 $<$0.01 6.12 8.22 0.10 0.09 0.05 0.29 13.1 0.8
Darknet 9 40 8.0 305.4 $<$0.01 $<$0.01 5.44 56.51 0.34 $<$0.01 0.03 $<$0.01 7.8 0.5
Software 14 28 5.4 40.1 $<$0.01 $<$0.01 6.50 9.61 0.42 0.42 0.10 0.36 17.4 1.4
Video&Films 21 23 9.6 61.1 $<$0.01 0.02 7.61 15.10 1.15 0.58 0.17 0.10 48.6 5.1
Politics 11 34 3.6 160.8 $<$0.01 0.07 9.69 22.17 3.03 0.25 0.17 0.26 21.2 1.9
Erotic 15 14 6.6 60.1 $<$0.01 $<$0.01 6.05 17.29 0.80 0.03 0.02 $<$0.01 60.1 2.8
Linguistics 20 11 12.6 35.0 0.10 0.08 2.34 15.34 0.33 0.16 $<$0.01 0.04 19.5 3.4

### v2

@@ 7. Ethical aspects
$\bullet$ Our data extraction rate on TGStat is below 5 pages per minute, to minimise the load, and we repeat the crawling only once a week.
@@
$\bullet$ We contacted Telegram’s support to declare our intention, asking them to share with us any limitation. We received no answer. The privacy policy of the platform does not forbid crawling.
@@
$\bullet$ To respect privacy restrictions imposed by group administrators, we restrict our analysis to public groups.We do not monitor groups where the administrator sets the auto-delete functionality and those where admins refuse or ignore our join request.
@@
$\bullet$ Telegram might be used to share copyright-protected material and illicit content. We avoid storing any non-textual element, only collecting metadata (e.g., video duration and size).
Acknowledgements
This work has been partially supported by the Spoke 1 ”FutureHPC & BigData” of ICSC — Centro Nazionale di Ricerca in High-Performance-Computing, Big Data and Quantum Computing, funded by European Union — NextGenerationEU. It has also been partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação de Amparo à Pesquisa do Estado de Minas Gerais (FAPEMIG), both in Brazil.
Appendix A Appendix
@@
Appendix includes additional results to contextualize our analysis.
In Figure 10, we show the breakdown of the groups we found on TGStat, and those we actually monitor for the paper (blue pattern). The figure details the fraction of groups we ignored for various motivations.
In Table 3, we provide a qualitative explanation of the topics under observation. In
Table 4, we present additional per-topic detailed statistics.
In Figure 4.1.1, we compare the median use of non-textual media in group messages by topic.
Figure 10. Per-topic groups of TGStat and results of our crawling.
Table 4. Per-topic additional metrics.
Topic English groups Russian groups Bot msgs (%) Avg msg length (ch.) Avg video size (MB) Avg video duration (min)
Education 46 13 8.9 43.3 49.9 4.3
Bookmaking 24 49 6.4 68.3 14.1 0.8
Crypto 45 15 8.3 44.4 12.5 0.9
Technologies 23 30 9.9 58.5 13.1 0.8
Darknet 9 40 8.0 305.4 7.8 0.5
Software 14 28 5.4 40.1 17.4 1.4
Video&Films 21 23 9.6 61.1 48.6 5.1
Politics 11 34 3.6 160.8 21.2 1.9
Erotic 15 14 6.6 60.1 60.1 2.8
Linguistics 20 11 12.6 35.0 19.5 3.4
All Topics 228 257 7.8 65.5 26.1 2.1